OpenStudy (anonymous):

Which numbers should be multiplied to obtain 183^2 - 139^2

OpenStudy (aric200):

14168?

OpenStudy (loser66):

it forms the form of a^2 -b^2 = (a+b)(a-b) so that, you just (add them) * (subtract them)

OpenStudy (loser66):

got me?

OpenStudy (loser66):

(183+139)(183-139) = simplify it to get the answer

OpenStudy (aric200):

which then = yeah

OpenStudy (aric200):

@BasedGod1122